Construct a Pareto chart for the number of unemployed. People in the selected States for June of 1995.
(States)=Taxes, New York, Pennsylvania, Florida, Ohio.
(Number)= 605,000 , 494,000 , 364,000 , 363,000 , 269,000
1
Expert's answer
2020-03-22T17:02:54-0400

Arrange the data from the largest to smallest according to frequency.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 State & Number \\ \hline
 Texas & 605000 \\
 \hdashline
 New York & 494000 \\ \hline
 Pennsylvania & 364000 \\ \hline
 Florida & 363000 \\ \hline
 Ohio & 269000 
\end{array}"

Draw and label the x and y axes. 

Draw the bars corresponding to the frequencies.

The Pareto chart shows that assaults had the highest frequency and homicides the lowest.

"Total=605000+494000+364000+363000+269000="

"=2095000"

"{605000 \over 2095000}\cdot100\%=28.878\%"

"{494000 \over 2095000}\cdot100\%=23.580\%""\ 28.878\%+23.580\%=52.458\%"

"{364000 \over 2095000}\cdot100\%=17.375\%""52.458\%+17.375\%=69.833\%"

"{363000 \over 2095000}\cdot100\%=17.327\%""69.833\%+17.327\%=87.160\%"

"{269000 \over 2095000}\cdot100\%=12.840\%""87.160\%+12.840\%=100\%"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 State & Number & Cum. Relative \ Frequencies (\%)\\ \hline
 Texas & 605000 & 28.88 \\
 \hdashline
 New York & 494000 & 52.46 \\ \hline
 Pennsylvania & 364000 & 69.83 \\ \hline
 Florida & 363000 & 87.16 \\ \hline
 Ohio & 269000 & 100 \\ \hline
 Total= & 2095000 & 
\end{array}"

Therefore, the following Pareto Chart is obtained based on the table above: